OpenStudy (ladiesman217):

Medals & Fans awarded 1.) Create a trinomial. Write it in ascending order and identify its degree. 2.) Part 1: Explain, in complete sentences, whether or not the polynomial x2 – 16x + 64 fits one of the special patterns. If so, which one? Part 2: Explain how the polynomial is factored. Part 3: Provide the factors. 3.) Explain, in complete sentences, how to factor 5x2 – 5x – 30. Explain how your factors could be checked for accuracy.

OpenStudy (ddcamp):

The main special patterns: (x+a)² = x² + 2ax + a² (x+a)(x-a) = x² - a²

OpenStudy (ladiesman217):

huh

OpenStudy (ddcamp):

For part 1, a trinomial is just a polynomial with 3 parts: If your trinomial is: c + bx + ax² The degree is 2 (x²) The degree is the largest power used.

OpenStudy (ladiesman217):

ok 2 & 3

OpenStudy (ddcamp):

2: Is the question in the form x² + 2ax + a² ? If so, it factors to (x+a)² x² - a² ? If so, it factors to (x+a)(x-a) (Difference of squares)

OpenStudy (ladiesman217):

ok

OpenStudy (ddcamp):

For 3: Divide each part by 5, leaving you with: x² - x - 6 (x+a)(x+b) a + b = -1 ab = -6